It takes a girl 45 minute to deliver the papers on her route; however if her brother helps, it takes only 20 minutes. How would it take her brother to deliver the paper by himself?
Answer by mananth(16949)

let delivering the papers be = 1 job
..
girl does it in 45 minutes
1/45 of the job she doe in 1 minute
..
together they do the job in 20 minutes
1/20 of the job they do in 1 minute.
..
1/20 -1/45 = job done by brother in 1 minute
1/20-1/45=(9-4 )/ 180
=5/180
=1/36 of the job he does in 1 minute
he will take 36 minutes to do the job.